If the present ages of son and father are expressed by x and y respectively and after ten years father will be twice as old as his son. Write the relation between x and y
Answer by mananth(16946)

son --x
father--------y
after 10 years
son x+10
father---y+10
y+10=2(x+10)
y+10=2x+20
y=2x+10